# IJCAI–25 Formatting Instructions

### v1

@@ 7 Conclusions and Outlook
In this paper, we have explored various approaches to take advantage of monolingual examples in retrieval-augmented neural machine translation. Our main conclusion, resulting from experiments on 2 language directions and 16 varied textual domains, is that retrieving directly in the target language can be as effective as fuzzy-matching on the source-side, especially with lexical-based fine-tuning objectives. Using these techniques, we were also able to obtain large BLEU increases over strong baselines on the Wikipedia dataset (up to +3.8 BLEU points for our best configuration), showcasing the benefits of RANMT in a large-scale setup. In our implementation, cross-lingual retrieval is no more costly than performing fuzzy matches in the source language.
We identify several ways to improve these results: (a) by fine-tuning the retrieval models on very large multilingual datasets, adopting data settings that are comparable to the pre-training of LASER/LaBSE; (b) using more recent sentence embedders such as SONAR (Duquenne et al., 2023); (c) using CLIR techniques to train the RANMT models, instead of relying on source-side matches as we have done here; (d) revisiting the other aspect of the retrieval pipeline – e.g., relaxing the filtering threshold during example selection. Another continuation of this work is to adopt CLIR for in-context learning with LLMs, which are currently instruction-tuned with parallel examples, based on source-side similarities Moslem et al. (2023). Fine-tuning LLMs with just target side examples retrieved with CLIR is one priority of our future work. Admittedly, we worked with high ressource languages to assess the viability of our method. Considering a low ressource setup, provided large enough monolingual corpora are available, is a natural use case when it is even harder to access to parallel data. Such experiments can further assess the strenght of our method.
@@ Appendix A Parallel Data Specification
The en-fr data is a clean version [^8] We filter noisy parallel pairs with COMETKiwi Rei et al. [2022] and prepare a new train/valid/test split. of the one used by [Xu et al., 2020]. Validation and test sets each contain $1,000$ segments. The de-en corpus is borrowed unchanged from [Aharoni and Goldberg, 2020] and is also used in [Cai et al., 2021; Agrawal et al., 2023; Bouthors et al., 2024]. The associated validation and test sets contain $2,000$ segments.
@@ Appendix B Search Parameterization
For both methods, we search up to $k$=3 examples.
@@ B.1 Fuzzy Matching
The execution of the fuzzy matching methods (fuzzy-src, fuzzy-bt and fuzzy-gold) is done via an open source library [^9] https://github.com/SYSTRAN/fuzzy-match. During the evaluation, there is no filter applied before scoring the segments with $\operatorname{Lev}$. However, the examples associated to the training samples are selected with $\operatorname{Lev}$and a BM25 filter. This ensures a reasonable time to build the set of examples for traning.
@@ B.2 FAISS
The execution of the CLIR techniques is done with library FAISS [Douze et al., 2024]. The $k$NN search is performed on GPU (IndexFlatIP), without IVF (for faster search) or quantizer (for lower memory footprint) to ensure the retrieval of the best matches. Indeed, both optimizations can prevent the model from finding the optimal $k$ nearest neighbors.
@@ Appendix C NMT Architectures Configuration
Both TM3-LevT and NFA models rely on the default backbone Transformer architecture [Vaswani et al., 2017]. We train one instance of TM3-LevT for each language pair (en-fr and de-en), using only the training data available in the corresponding datasets. Our NFA models are trained on much larger sets of parallel data, including a good share of Web data. We use these models to showcase the effectiveness of cross-lingual retrieval for RAMNT systems trained with large, high-quality, parallel datasets (8M sentences for en-fr and 11.5M for de-en). After inspecting the training datasets of NFA models, we found that some of them partly overlap with the training data; the corresponding contamination rates are reported in Table 4.
@@
In addition, we use the multilingual NLLB model [Costa-jussà et al., 2024] out-of-the-box for both language pairs.101010version: NLLB-200-distilled-1.3B
@@
Appendix D Computational Cost of Retrieval
@@
The computational cost can be dived into two categories: a fixed cost (independant from the number of sentences to translate) and a variable cost (proportional to this number). The retrieval process for training falls into the former, as well as the encoding of the whole set of monolingual target segments (for CLIR techniques) and the indexation of the TM (for fuzzy matching techniques). The variable costs lie in the fuzzy match search with filtering and the computation of $\operatorname{Lev}$, and in the CLIR search with the encoding of the source sentences, then the $k$NN search. In our experiments, we removed the filter (BM25) so that we obtain the best possible fuzzy matches (fuzzy-src, fuzzy-bt and fuzzy-gold), at the cost of a $\sim$100 times higher latency. However, when applying a BM25 filter to preselect 100 segments, we observe that both fuzzy matching and CLIR techniques can retrieve their closest match in about $\sim$1ms. We obtained this result un corpus ECB (150K sentences) with optimmal conditions:
@@
  • as for CLIR, source sentences are encoded in batches of size 50, and the $k$NN search is handled by library FAISS. We used a V100-32GB GPU.

### v2

@@ 7 Conclusions and Outlook
In this paper, we have explored various approaches to take advantage of monolingual examples in retrieval-augmented neural machine translation. Our main conclusion, resulting from experiments on 3 language directions and 16 varied textual domains, is that retrieving directly in the target language can be as effective as fuzzy-matching on the source-side, especially with lexical-based fine-tuning objectives. Using these techniques, we were also able to obtain large BLEU increases over strong baselines on the Wikipedia dataset (up to +3.8 BLEU points for our best configuration), showcasing the benefits of RANMT in a large-scale setup. In our implementation, cross-lingual retrieval is no more costly than performing fuzzy matches in the source language.
We identify several ways to improve these results: (a) by fine-tuning the retrieval models on very large multilingual datasets, adopting data settings that are comparable to the pre-training of LASER/LaBSE; (b) using more recent sentence embedders such as SONAR (Duquenne et al., 2023); (c) using CLIR techniques to train the RANMT models, instead of relying on source-side matches as we have done here; (d) revisiting the other aspect of the retrieval pipeline – e.g., relaxing the filtering threshold during example selection.
Limitations
In this work, we have focused on extending a translation setting that is common in the translation industry, where translation memories have repeatedly been shown to speed up the translation process and improve the consistency of the resulting texts. In such setting, it is also common to have access to large monolingual resources, as has been studied eg., by (Tezcan et al., 2024) for the legal domain and for (Tamura et al., 2023) in the scientific domain. By design, this setting is only applicable for so-called “high-resource" language pairs, in which high-quality systems can be built and improved upon using auxiliary resources. In our selection of language pairs, we have favored domain diversity over language diversity, to better highlight how the availability of relevant examples impacts the overall translation quality. As our experiments show, the effectiveness of the cross-lingual approach does not depend on the similarity of the source and target languages, but mostly relies on the quality of cross-lingual sentence alignment. As shown e.g., by (Feng et al., 2022), very precise alignments can be achieved, even for distant language pairs such as English-Chinese.
We have thus deliberately chosen to exclude translation from and into “low-resource” languages, for which many resources and components are still difficult to collect (eg., parallel corpora to train sufficiently good NMT baselines, to be used for back translation; and also training and test data in specialized domains, where the TM-based approach is most likely to retrieve very good matches). We reckon that also improving generic NMT systems that could handle translations for those language pairs is another important research area, albeit distinct from what we have chosen to study in this work.
Ethical Statement
There are no ethical issues.
Acknowledgments
This research was funded by the French Agence Nationale de la Recherche (ANR) under the project TraLaLaM (“ANR-23-IAS1-0006”). This work was performed using HPC resources from GENCI–IDRIS (Grant 2024-A0161015117).
@@ Appendix A Parallel Data Specification
The en-fr data is a clean version [^12] We filter noisy parallel pairs with COMETKiwi Rei et al. (2022) and prepare a new train/valid/test split. of the one used by (Xu et al., 2020). Validation and test sets each contain $1,000$ segments. The de-en corpus is borrowed unchanged from (Aharoni and Goldberg, 2020) and is also used in (Cai et al., 2021; Agrawal et al., 2023; Bouthors et al., 2024). The associated validation and test sets contain $2,000$ segments. The per-domain corpus sizes are available in table 8. The en-uk corpus is the one provided by (Tezcan et al., 2024), containing about 286K training sentences, 2000 validation sentences and 1898 test sentences. It corresponds to various legal texts.
Note that the preprocessing stages were performed with subword-nmt (Sennrich et al., 2016b) for tokenization and BPE, and fairseq (Ott et al., 2019) for binarization.
@@ Appendix B Search Parameterization
For both methods, we search up to kk=3 examples.
@@ B.1 Fuzzy Matching
The execution of the fuzzy matching methods (fuzzy-src, fuzzy-bt and fuzzy-gold) is performed using an open source library [^13] https://github.com/SYSTRAN/fuzzy-match. During the evaluation, there is no filter applied before scoring the segments with $\operatorname{Lev}$. However, the examples associated to the training samples are selected with $\operatorname{Lev}$and a BM25 filter. This ensures a reasonable time to build the set of examples for training.
@@ B.2 FAISS
The execution of the CLIR techniques is done with library FAISS (Douze et al., 2024). The kkNN search is performed on GPU (IndexFlatIP), without IVF (for faster search) or quantizer (for lower memory footprint) to ensure the retrieval of the best matches. Indeed, both optimizations can prevent the model from finding the optimal $k$ nearest neighbors.
@@ Appendix C NMT Architectures Configuration
Both TM3-LevT and NFA models rely on the default backbone Transformer architecture (Vaswani et al., 2017). We train one instance of TM3-LevT for each language pair (en-fr and de-en), using only the training data available in the corresponding datasets. Our NFA models are trained on much larger sets of parallel data, including a good share of Web data. We use these models to showcase the effectiveness of cross-lingual retrieval for RAMNT systems trained with large, high-quality, parallel datasets (8M sentences for en-fr and 11.5M for de-en). After inspecting the training datasets of NFA models, we found that some of them partly overlap with the training data; the corresponding contamination rates are reported in Table 4.
@@
In addition, we use the multilingual NLLB model (Costa-jussà et al., 2024) out-of-the-box for both language pairs.141414version: NLLB-200-distilled-1.3B
@@
Appendix D Adapting EuroLLM
LLMs are well suited for in-context learning in $k$-shot settings (Radford et al., 2019): a set of $k$ demonstrations of the task are presented to the model so that the completion of the task fulfills the prompt pattern/template. In our case, the situation is slightly different as we only include target-side examples, which are not suitable demonstrations of the translation task.
To adapt an LLM to this setting, we first compare (a) prompting and (b) fine-tuning strategies on the Wikipedia domain. [^15] Another alternative, that we do not consider in this work, would be to use backtranslation to automatically craft the missing source for in-context examples.
Regarding (a), we try to take advantage of the in-context abilities of the Instruct version of EuroLLM-9B, simulating in-context examples with the follow prompt patterns:
  • naive:
    Consider the following French similar translations.
    {k examples in French}
    Translate the following text from English to French.
    English:
    {source text}
    French:
  • empty:
    Consider the following English-French examples.
    English:
    French:
    {French example #1}
    Translate the following text from English to French.
    English:
    {source}
    French:
  • copy:
    Consider the following English-French examples.
    English:
    {French example #1}
    French:
    {French example #1}
    Translate the following text from English to French.
    English:
    {source}
    French:
The naive prompt gives a simple description of the task, where only the target side of examples are introduced. The copy and empty prompts simulate a standard $k$-shot prompt, where the missing source text is either left empty or contains a copy of the target text. We compare these prompts with a 0-shot setting, where no example is provided to the model (starting directly with "Translate the following text" for any of the three templates).
Regarding (b), we fine-tune the base EuroLLM-9B model with the naive prompt with up to 20K sentences from the training set for each domain. Fine-tuning is performed with QLoRA (Dettmers et al., 2023) with a rr=16, $\alpha$=32, dropout=0.05, no bias, applied on all linear layers, and 4-bit quantization. Training is performed with huggingface implementation [^16] https://huggingface.co/docs/transformers/trainer with a batch size of 32 for a single epoch and a learning rate of 2e-10. The in-context examples used for training are retrieved using the fuzzy-src strategy.
Instruct ft-base
0-shot naive copy empty naive
(k=0) (k=3) (k=3) (k=3) (k=3)
36.5 36.5 38.4 37.5 49.5
Table 9: BLEU score of various EuroLLM-9B based systems on the Wikipedia test set (0-shot and 33-shot). Target examples are retrieved with LaBSE with a custom threshold of $0.6$.
We observe the superiority of the fine-tuning strategy compared to the various prompt-based approaches in Table 9. As some of these differences might be explained by the implicit language and domain adaptation that takes place during model fine-tuning, we additionally fine-tune the base model with a standard $k$-shot prompt, including the source and target side of each examples. This setting directly compares with the instruct model, the only difference being the adaptation that is performed during fine-tuning; it also allows us to assess the performance loss that happens when using only target side examples. Results are in Table 10 for a mixture of $5$ domains.
EuroLLM model source k BLEU
Instruct - 0 42.9
3 54.0
ft-base (src+tgt) - 0 46.7
3 58.5
ft-base (tgt) - 0 47.3
3 57.5
Table 10: Average BLEU score of various EuroLLM-9B system fine-tuning schemes for ECB, EMEA, Europarl, JRC-Acquis and Wikipedia test sets (k-shot). Parallel examples are retrieved with fuzzy-src using a custom threshold of 0.
In Table 10, we observe that in fact the domain/language adaptation of EuroLLM accounts for about +4 BLEU points (ft-base (src+tgt) vs. instruct). For the sole Wikipedia domain, this adapation yields about +6 BLEU points improvement. This confirms that method (b) is in fact clearly superior to simple prompting strategies. Therefore, this is the method used for all the results reported in Section 6. Moreover, we observe that the absence of the source side of $k$-shot examples is slightly detrimental to the model performance ($\approx$ -1 BLEU point between the two fine-tuned versions, both for the 0-shot and $3$-shot settings), suggesting that the target side of the retrieved examples contains most of the information needed to guide the model towards the desired translation.
Appendix E Computational Cost of Retrieval
@@
The computational cost can be dived into two categories: a fixed cost (independant from the number of sentences to translate) and a variable cost (proportional to this number). The retrieval process for training falls into the former, as well as the encoding of the whole set of monolingual target segments (for CLIR techniques) and the indexation of the TM (for fuzzy matching techniques). The variable costs lie in the fuzzy match search with filtering and the computation of $\operatorname{Lev}$, and in the CLIR search with the encoding of the source sentences, then the kkNN search. In our experiments, we removed the filter (BM25) so that we obtain the best possible fuzzy matches (fuzzy-src, fuzzy-bt and fuzzy-gold), at the cost of a $\sim$100 times higher latency. However, when applying a BM25 filter to preselect 100 segments, we observe that both fuzzy matching and CLIR techniques can retrieve their closest match in about $\sim$1ms. We obtained this result un corpus ECB (150K sentences) with optimmal conditions:
@@
  • as for CLIR, source sentences are encoded in batches of size 50, and the kkNN search is handled by the FAISS library. We used a V100-32GB GPU.
Appendix F Retrieval and Translation Examples
retriever $\operatorname{Lev}$
reference Le mannequin apparaît dans plusieurs séries télévisées au début des années 2000. (The model also appears in several TV shows in the early 2000s.)
train LaBSE 0.14 Cette série fut diffusée sur TF1 pendant l’été 2000. (This show was broadcast on TF1 during the summer of 2000.)
train ft-LaBSE-MAE 0.23 Elle est aussi apparue dans des séries télévisées. (She also played in some TV shows.)
mono LaBSE 0.46 Elle se fait connaître par sa série au début des années 2000. (She made a name for herself thanks to her show in the early 2000s.)
mono ft-LaBSE-MAE 0.62 Elle débute dans des séries télévisées au début des années 1990. (She made her debut in some TV Shows in the early 1990s.)
reference Shasta a connu une histoire explosive et éruptive. (Shasta has had an explosive and eruptive history.)
train LaBSE 0.06 L’histoire de la région du Huaynaputina est marquée par un magma riche en silice. (The history of the Huaynaputina region is marked by silica-rich magma. )
train ft-LaBSE-MAE 0.44 Elle a une histoire très singulière. (She has a very peculiar history.)
mono LaBSE 0.07 Il s’agit d’un des cônes volcaniques du mont Shasta. (This is one one the vulcanic cones of Mount Shasta.)
mono ft-LaBSE-MAE 0.56 Elle a connu une histoire riche en rebondissements. (Its history is full of twists and turns.)
Table 11: Two illustrations of retrieved examples for the Wikipedia test set w.r.t. the chosen cross-lingual retriever, and the retrieval pool (train set vs. monolingual corpus). Exact matches are marked in bold. Indicative translations of each French segment into English are also provided.
An illustration of the retrieved examples is provided is table 11. It highlights the enhanced capability of ft-LaBSE-MAE to retrieve lexically closer examples, rather than semantically close ones. This is particularly clear in the second example, which mentions Shasta, a volcano. While LaBSE retrieves topic-related segments, ft-LaBSE-MAE tends to select off-topic sentences that nonetheless contain a higher number of shared lexical items.
Appendix G Back-Translation
Back-translation (BT) has long been identified as a very effective technique to handle monolingual data in Statistical Machine-Translation (Bojar and Tamchyna, 2011), then in Neural Machine Translation (Sennrich et al., 2016a; Currey et al., 2017; Edunov et al., 2018; Burlot and Yvon, 2018; Caswell et al., 2019). Assuming the task is to translate from L1 into L2, and that monolingual texts in L2 are available, the BT approach translates these “backward” into L1, to obtain an artificial parallel corpus, pairing automatically generated source sentences with actual human-written target sentences. In NMT, it is custom to jointly use “natural” and “artificial” subsets of data to train and/or adapt translation models. This is illustrated in the first two columns (a) and (b) in Table 12, using no-retrieval ($k=0$), and corresponding respectively to the default (only parallel data) and BT-augmented settings. [^17] We omit the case where only artificial data is used, as in fully unsupervised machine translation (Lample et al., 2018).
MT setting (a) (b) (c) (d) (e) (f)
Parallel data
train ($k=0$) - - - -
train ($k\geq 0$) - -
test ($k\geq 0$) - -
BT-ed data
train ($k=0$) - - - - -
train ($k\geq 0$) - - - -
test ($k\geq 0$) - - - -
Table 12: Using back-translation in NMT and RANMT
Column (c) in Table 12 illustrates standard RANMT, where $k\geq 0$ parallel examples are used during training and at inference; columns (d), (e) and (f) correspond to three possible ways to integrate artificial data: (d) only during training, and retrieve selectively from the high-quality parallel data; (e) only during inference, extending a generic RANMT with relevant data; (f) both during training and at inference as in Tezcan et al. (2024). Our use of back-translation in the experiments of Section 5 corresponds to column (e), a setting we deem representative of typical use of monolingual data in actual applications, and which crucially does not require retraining stage. Note that the same scenario is used in our CLIR approach, which, in addition, fully dispenses with the back-translation stage.
Appendix H Experiments in English-Ukrainian
To further emphasize the ability of our method to retrieve relevant examples even in the absence of direct lexical overlap between source and target segments, we experiment with the translation from English into Ukrainian, with the source using the Latin script, and the target the Cyrillic script. We used the same dataset as Tezcan et al. (2024), corresponding to legal texts. The parallel corpus is split into train, validation and test sets, each containing respectively 286,417, 2000 and 1898 sentences. This corpus is augmented with a monolingual in-domain dataset of 1,461,320 Ukrainian sentences.
Our experiments are conducted with the same NFA architecture as the experiments in Section 5, as it produces translations with higher estimated-quality than the non-autoregressive system. This NFA model has been trained on 5M en-uk parallel pairs from various domains, including the in-domain train set.
Retrieval method Corpus BLEU COMET
no example - 51.2 92.7
fuzzy-src train 54.2 92.8
LaBSE train 53.9 92.7
ft-LaBSE-MSE train 54.3 92.8
LaBSE train+mono 57.2 92.9
ft-LaBSE-MSE train+mono 57.7 93.0
Table 13: BLEU and COMET scores obtained by the NFA model according to the retrieval method. Significance w.r.t. to the LaBSE counterpart is indicated by .
We fine-tune LaBSE using the proposed MSE lexical loss, with examples retrieved via fuzzy-src from the training set. We then evaluate and compare the performance of fuzzy-src, LaBSE, and ft-LaBSE-MSE by computing BLEU and COMET scores of translations produced by the NFA model. For LaBSE and ft-LaBSE-MSE, we consider two retrieval settings — train and train+mono — which correspond to the datasets from which target-side examples are retrieved for the CLIR approach. Results are in Table 13, scores are computed respectively by SacreBLEU and COMET-22, significance tests also use the SacreBLEU implementations with paired bootstrap resampling (n=1000, p=0.05).
We can make the following observations: (a) the baseline NFA model is about 3 BLEU points better than baseline encoder-decoder model; (b) CLIR here again matches the default RAMT set-up, where retrieval only exploits the parallel training data; (c) augmenting the retrieval pool with monolingual data again induces a significant performance boost (+3.4 BLEU), when retrieval relies on the fine-tuned version of LaBSE. These observations are consistent with what is reported in the main text.
Appendix I COMET Scores
For the full picture, we also report the COMET scores of each domain for all our experiments in Tables 14, 15 and 16.
English-French German-English
ECB EME Epp GNO JRC KDE News PHP TED Ubu Wiki KDE Kor JRC EME Sub
fuzzy-gold 70.5 85.0 45.8 73.2 71.9 58.9 24.3 35.9 21.2 52.1 29.8 30.3 -98.3 19.9 29.7 -24.8
fuzzy-src 69.0 81.5 44.7 67.6 69.8 47.9 24.7 33.5 19.5 49.2 25.0 24.1 -85.8 16.6 21.8 -28.2
fuzzy-bt 61.6 75.0 45.6 55.5 65.8 37.3 24.1 30.1 17.7 45.4 24.2 11.4 -110.6 -11.0 21.3 -25.4
LASER 65.8 76.3 45.1 58.1 66.0 34.7 23.2 29.2 19.9 45.0 24.5 18.5 -73.0 19.7 27.5 -31.0
LaBSE 65.5 77.2 42.3 62.9 67.9 42.8 23.3 29.7 19.0 47.1 23.3 21.2 -86.4 21.4 26.3 -30.1
ft-LaBSE-MSE 68.0 81.1 44.4 66.5 69.8 48.0 24.4 31.9 18.5 49.6 23.8 18.5 -76.3 8.7 15.6 -29.1
ft-LaBSE-MAE 69.0 80.7 44.2 65.5 68.4 47.0 23.8 32.7 18.9 49.8 24.9 20.2 -70.6 14.1 19.1 -28.6
ft-LaBSE-Rank 67.3 79.6 44.8 63.4 68.8 46.5 23.8 32.2 19.7 47.9 24.5 20.7 -84.9 14.0 17.6 -28.5
Table 14: Per domain COMET scores for TM3-LevT.
English-French German-English
ECB EME Epp GNO JRC KDE News PHP TED Ubu Wiki KDE Kor JRC EME Sub
no example 58.9 55.9 39.9 48.8 62.3 42.8 50.1 45.6 43.9 43.5 36.4 82.2 68.9 83.2 84.7 80.2
fuzzy-gold 90.5 90.9 87.9 89.8 90.7 85.8 88.2 84.6 85.4 86.6 85.2 84.5 71.4 88.2 86.3 80.6
fuzzy-src 90.4 90.5 87.8 89.3 90.5 85.2 88.2 84.5 85.4 86.3 85.0 84.1 70.0 87.9 85.9 80.4
fuzzy-bt 90.0 90.1 87.7 88.6 89.9 84.4 88.2 84.3 85.2 86.0 84.9 83.5 70.1 87.5 85.2 80.3
LASER 90.0 90.1 87.8 88.8 90.0 84.4 88.2 84.2 85.4 86.4 84.7 83.5 70.0 87.4 85.3 80.4
LaBSE 90.0 90.3 87.5 89.0 89.8 84.7 88.2 84.1 85.3 86.4 84.9 83.8 70.3 87.2 85.5 80.3
ft-LaBSE-MSE 90.3 90.5 87.7 89.2 90.4 85.1 88.2 84.5 85.3 86.5 84.8 83.8 70.3 87.8 85.6 80.4
ft-LaBSE-MAE 90.3 90.5 87.7 89.2 90.4 85.1 88.2 84.5 85.4 86.4 84.8 83.7 70.2 87.7 85.7 80.3
ft-LaBSE-Rank 90.2 90.4 87.7 89.2 90.5 85.3 88.2 84.5 85.3 86.5 84.9 83.8 70.5 87.8 85.7 80.4
Table 15: Per domain COMET scores for NFA.
English-French German-English
ECB EME Epp GNO JRC KDE News PHP TED Ubu Wiki KDE Kor JRC EME Sub
no example 88.2 88.5 87.7 85.3 89.1 82.8 86.2 84.3 85.0 85.8 86.3 85.9 74.7 86.9 85.1 80.5
fuzzy-gold 90.2 90.9 87.7 90.1 90.3 87.5 86.3 85.2 85.4 89.0 87.0 88.0 75.5 88.4 87.0 80.8
fuzzy-src 90.0 90.5 87.8 89.3 90.1 86.2 86.2 85.1 85.3 88.4 86.8 86.5 74.9 87.2 85.5 80.7
fuzzy-bt 88.6 89.0 87.7 87.2 89.2 84.0 86.2 84.7 85.3 87.6 86.6 86.0 74.9 87.2 85.6 80.6
LASER 89.8 90.0 87.7 88.1 90.0 84.6 86.3 84.9 85.2 88.3 86.6 86.3 75.0 87.9 85.9 80.6
LaBSE 90.0 90.2 87.7 88.6 90.1 85.8 86.3 84.9 85.2 88.2 86.6 86.5 75.0 87.9 86.1 80.5
ft-LaBSE-MSE 90.0 90.5 87.7 88.9 90.1 86.4 86.3 85.1 85.3 88.5 86.8 86.5 75.1 87.8 85.8 80.6
ft-LaBSE-MAE 89.9 90.6 87.7 89.1 90.1 86.4 86.2 85.0 85.2 88.4 86.7 86.7 75.1 87.9 86.0 80.5
ft-LaBSE-Rank 89.8 90.3 87.7 89.2 90.0 86.2 86.3 84.9 85.3 88.6 86.6 86.6 75.0 87.8 86.0 80.5
Table 16: Per domain COMET scores for EuroLLM.